### v2

# Semi-Riemannian $\text{spin}^{c}$ manifolds carrying generalized Killing spinors and the classification of Riemannian $\text{spin}^{c}$ manifolds admitting a type I imaginary generalized Killing spinor

Abstract
We classify Riemannian $\text{spin}^{c}$ manifolds carrying a type I imaginary generalized Killing spinor, by explicitly constructing a parallel spinor on each leaf of the canonical foliation given by the Dirac current. We also provide a class of Riemannian $\text{spin}^{c}$ manifolds carrying a type II imaginary generalized Killing spinor, by considering spacelike hypersurfaces of Lorentzian $\text{spin}^{c}$ manifolds. We carry out much of the work in the setting of semi-Riemannian $\text{spin}^{c}$-manifolds carrying generalized Killing spinors, allowing us to draw conclusions in this setting as well. In this context, the Dirac current is not always a closed vector field. We circumvent this in even dimensions, by considering a modified Dirac current, which is closed in the cases when the original Dirac current is not. On the path to these results, we also study semi-Riemannian manifolds carrying closed and conformal vector fields.
keywords:
$\text{spin}^{c}$ geometry, generalized Killing spinors, Dirac current, closed & conformal vector fields
1 Introduction
We say that a non-trivial (complex) spinor field $\varphi$ on a semi-Riemannian $\text{spin}^{c}$ manifold $(M,g)$ is a generalized Killing spinor if there is some $g$-symmetric endomorphism field $A$ of $TM$ such that
$\nabla^{S}_{X}\varphi=jA(X)\cdot\varphi\hskip 14.22636pt\forall X\in\Gamma(TM),$ (1)
where $j\in\{1,i\}$. We call $A$ the Killing endomorphism and if $j=1$ $(j=i)$ we say that $\varphi$ is a real (imaginary) generalized Killing spinor. This is not the original definition from [37] of a generalized Killing spinor, but it is the one that is most frequent. The original definition only allowed for the Killing endomorphism $A$ to be of the form $A=\mu\operatorname{Id}$, where $\mu$ is some real valued function on $M$. For this reason, we introduce the term ”special generalized Killing spinor” to refer to this case. When the Killing endomorphism $A$ is a constant multiple of the identity, we simply call the spinor field a Killing spinor.
Killing spinors where first considered in physics. They were originally introduced in general relativity and later appeared in supergravity theories. See the introduction of [9] for a full description of their origin. Killing spinors are also of purely geometric interest. The existence of a Killing spinor on a semi-Riemannian spin manifold implies that the manifold has constant scalar curvature, and in the Riemannian case, the manifold has to be Einstein [9]. Real Killing spinors on compact Riemannian spin manifolds appear as the limiting case of Friedrich’s eigenvalue estimate for the Dirac operator [15]. Killing spinors are also related to supergeometry [1]. Generalized Killing spinors appear on hypersurfaces of semi-Riemannian spin and $\text{spin}^{c}$ manifolds carrying parallel spinors, which is a motivation for their study [6, 3, 13, 12].
Riemannian spin manifolds with parallel spinors are Ricci-flat and have special holonomy group [27, 16, 39]. Baum classifies complete Riemannian spin manifolds carrying imaginary Killing spinors in [8] and [7]. Baum splits the classification into two cases, called type I and type II imaginary Killing spinors. For the type I case, Baum finds a suitable foliation of the manifold consisting of hypersurfaces carrying parallel spinors. From such a hypersurface, it is shown that a flow in the normal direction gives rise to a warped product, isometric with the original manifold. For the type II case, Baum shows that the manifold must be isometric to hyperbolic space. The classification of complete Riemannian spin manifolds with real Killing spinors goes back to the group of Friedrich, and their collaborators [18, 19, 20, 21, 24]. The full classification of complete, simply connected Riemannian spin manifolds carrying real Killing spinors is due to Bär [4]. This is done by showing that the cone over the given manifold carries a parallel spinor.
The above results have been extended by many in various directions. Among these, the following classifications are directly related to this article. Moroianu [35] classifies complete, simply connected Riemannian $\text{spin}^{c}$ manifolds carrying parallel spinors. Moroianu shows that the manifold splits as a Riemannian product between a Kähler manifold (equipped with its canonical $\text{spin}^{c}$-structure) and a spin manifold carrying a parallel spinor [35, Theorem 1.1]. Shortly after Baum, Rademacher extends Baum’s results to what we call special imaginary generalized Killing spinors [37]. Große and Nakad have extended Rademacher’s classification to the $\text{spin}^{c}$ setting [23, Theorem 4.1]. An extension of Rademacher’s result to the setting of Riemannian spin manifolds carrying a type I imaginary generalized Killing spinor have been made by Leistner and Lischewski [33, Theorem 4].
2 Preliminaries and notation
This section is a summary of some preliminaries from spin and $\text{spin}^{c}$ geometry. The reader is directed to [31, 14, 17] for details.
For $n>1$, let $\theta:\text{Spin}_{\text{GL}}(n)\to\text{GL}_{n}^{+}$, be the connected double covering group of $\text{GL}_{n}^{+}$, where $\text{GL}_{n}^{+}$ is the connected component of the identity in $\text{GL}_{n}(\mathbb{R})$. If $n>2$, this corresponds to the universal covering group. Let $M$ be an oriented smooth manifold of dimension $n$ and let $P_{\text{GL}^{+}}$ be the bundle of oriented frames on $M$. A topological spin structure on $M$ is a $\text{Spin}_{\text{GL}}(n)$-principal bundle $P_{\text{Spin}_{\text{GL}}}\to M$, such that the associated $\text{GL}_{n}^{+}$-principal bundle $P_{\text{Spin}_{\text{GL}}}\times_{\theta}\text{GL}_{n}^{+}$ coincides with $P_{\text{GL}^{+}}$, meaning there is a bundle isomorphism $I:P_{\text{Spin}_{\text{GL}}}\times_{\theta}\text{GL}_{n}^{+}\to P_{\text{GL}^{+}}$, covering the identity. We say that two topological spin structures $P_{\text{Spin}_{\text{GL}}}$ and $P_{\text{Spin}_{\text{GL}}}^{\prime}$ are isomorphic if there is an isomorphism $P_{\text{Spin}_{\text{GL}}}\to P_{\text{Spin}_{\text{GL}}}^{\prime}$ of $\text{Spin}_{\text{GL}}(n)$-principal bundles, compatible with the corresponding isomorphisms $I:P_{\text{Spin}_{\text{GL}}}\times_{\theta}\text{GL}_{n}^{+}\to P_{\text{GL}^{+}}$ and $I^{\prime}:P_{\text{Spin}_{\text{GL}}}^{\prime}\times_{\theta}\text{GL}_{n}^{+}\to P_{\text{GL}^{+}}$. Further, we say that $\text{Spin}_{\text{GL}}^{c}(n):=\text{Spin}_{\text{GL}}(n)\times_{\mathbb{Z}/2}S^{1}$ is the topological $\text{spin}^{c}$ group and we note that there are two natural maps, $\lambda:\text{Spin}_{\text{GL}}^{c}(n)\to\text{GL}_{n}^{+}$ and $\rho:\text{Spin}_{\text{GL}}^{c}(n)\to S^{1}$, defined via $\lambda([A,z]):=\theta(A)$ and $\rho([A,z])=z^{2}$. We define a topological $\text{spin}^{c}$ structure on $M$ to be a $\text{Spin}_{\text{GL}}^{c}(n)$-principal bundle $P_{\text{Spin}_{\text{GL}}^{c}}\to M$ such that the associated $\text{GL}_{n}^{+}$-principal bundle $P_{\text{Spin}_{\text{GL}}^{c}}\times_{\lambda}\text{GL}_{n}^{+}$ coincides with $P_{\text{GL}^{+}}$. The isomorphism classes of such structures are defined in analogy with the aforementioned spin case. To any topological $\text{spin}^{c}$ structure, we have a naturally associated $S^{1}$-principal bundle $P_{S^{1}}:=P_{\text{Spin}_{\text{GL}}^{c}}\times_{\rho}S^{1}$. A topological spin structure $P_{\text{Spin}_{\text{GL}}}$ gives rise to a double covering $P_{\text{Spin}_{\text{GL}}}\to P_{\text{GL}^{+}}$ of principal bundles. Similarly, a topological $\text{spin}^{c}$ structure $P_{\text{Spin}_{\text{GL}}^{c}}$ gives rise to a double covering $\Lambda:P_{\text{Spin}_{\text{GL}}^{c}}\to P_{\text{GL}^{+}}\oplus P_{S^{1}}$ of principal bundles. Note that a topological spin structure $P_{\text{Spin}_{\text{GL}}}$ induces a topological $\text{spin}^{c}$ structure via $P_{\text{Spin}^{c}_{\text{GL}}}:=P_{\text{Spin}_{\text{GL}}}\times_{\mathbb{Z}/2\mathbb{Z}}P_{S^{1}}$, where $P_{S^{1}}$ denotes the trivialized $S^{1}$-principal bundle on $M$.
Remark 1.
In terms of classifying spaces, a topological $\text{spin}^{c}$ structure on a smooth manifold $M$ of dimension $n$ is a lift $M\to\text{BSpin}^{c}_{\text{GL}}(n)$ of the classifying map $M\to\text{BGL}^{+}_{n}$ for $P_{\text{GL}^{+}}$ along the natural map $\text{BSpin}^{c}_{\text{GL}}(n)\to\text{BGL}^{+}_{n}$ induced by the map $\lambda:\text{Spin}^{c}_{\text{GL}}(n)\to\text{GL}^{+}_{n}$ from above. An isomorphism class of topological $\text{spin}^{c}$ structures then corresponds to a homotopy class of such lifts. The corresponding remark for topological spin structures is analogous.
Next, let $g$ be a semi-Riemannian metric on $M$ of index $r$ such that $(M,g)$ is space and time oriented. We allow $r=0$, corresponding to the Riemannian case, but we exclude the case $r=n$. This yields a reduction $\iota:P_{\text{SO}^{0}}\xhookrightarrow{}P_{\text{GL}^{+}}$ of principal bundles along the canonical inclusion $\text{SO}^{0}(r,n-r)\xhookrightarrow{}\text{GL}^{+}$, where $\text{SO}^{0}(r,n-r)$ denotes the identity component of $\text{SO}(r,n-r)$. Given a topological $\text{spin}^{c}$ structure on $M$, we define the corresponding (geometric) $\text{spin}^{c}$ structure as the pullback $P_{\text{Spin}^{c}}$ along $\Lambda$ and $\iota\oplus\operatorname{Id}:P_{\text{SO}^{0}}\oplus P_{S^{1}}\to P_{\text{GL}^{+}}\oplus P_{S^{1}}$. In other words, the principal bundle $P_{\text{Spin}^{c}}$ fits into the following pullback diagram
$P_{\text{Spin}^{c}}$$P_{\text{Spin}_{\text{GL}}^{c}}$$P_{\text{SO}^{0}}\oplus P_{S^{1}}$$P_{\text{GL}^{+}}\oplus P_{S^{1}}.$}} $\scriptstyle{\Lambda}$$\scriptstyle{\iota\oplus\text{id}}$
The definition of a spin structure is analogous. A space and time oriented semi-Riemannian manifold $(M,g)$, equipped with a $\text{spin}^{c}$ (spin) structure is said to be a time oriented semi-Riemannian $\text{spin}^{c}$ (spin) manifold.
Let $(M,g)$ be a time oriented semi-Riemannian $\text{spin}^{c}$ manifold of dimension $n$ and index $r$. The bundle $P_{\text{spin}^{c}}$ is a $\text{spin}^{c}(r,n-r)$-principal bundle, where $\text{spin}^{c}(r,n-r):=\text{spin}(r,n-r)\times_{\mathbb{Z}/2\mathbb{Z}}S^{1}$ and where we write $\text{spin}(r,n-r)$ to denote the connected component of the identity of the spin group sitting inside the Clifford algebra $C\ell_{r,n-r}$. Letting $\delta:\text{spin}(r,n-r)\to\mathrm{GL}(\Sigma)$ be the restriction of an irreducible representation of the complexified Clifford algebra $\mathbb{C}\ell_{r,n-r}$, we define $\Delta:\text{spin}^{c}(r,n-r)\to\mathrm{GL}(\Sigma)$ via $\Delta([s,z])v=z\delta(s)v$. The associated vector bundle $S:=P_{\text{spin}^{c}}\times_{\Delta}\Sigma$ is called the complex spinor bundle $S\to M$, which is a vector bundle of complex dimension $2^{\lfloor\frac{n}{2}\rfloor}$. When $n$ is even, $S$ splits as a direct sum $S=S^{+}\oplus S^{-}$.
Structures on the spinor bundle.
In any dimension, the spinor bundle $S$ carries the following structures:
  1. 1.
    Clifford multiplication $\cdot:TM\otimes S\to S$, $X\otimes\varphi\mapsto X\cdot\varphi$, which satisfies the Clifford relation
    $X\cdot Y\cdot\varphi+Y\cdot X\cdot\varphi=-2g(X,Y)\varphi.$
    Using the metric, we can also define Clifford multiplication by one-forms. In case $n$ is even, Clifford multiplication maps elements $\varphi\in S^{+}$ to $X\cdot\varphi\in S^{-}$.
  2. 2.
    A hermitian bundle metric $\langle{\cdot,\cdot}\rangle$, which is positive definite when $r=0$ and has signature $2^{\lfloor\frac{n}{2}\rfloor-1}$ when $r>0$. It obeys the rule
    $\langle{X\cdot\varphi,\psi}\rangle=(-1)^{r+1}\langle{\varphi,X\cdot\psi}\rangle.$
    In case $n$ is even, the bundle metric makes the splitting $S^{+}\oplus S^{-}$ orthogonal.
  3. 3.
    Given a connection $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$, we obtain a metric connection $\nabla^{S}$ on $S$ which satisfies
    $\nabla_{X}^{S}(Y\cdot\varphi)=(\nabla_{X}Y)\cdot\varphi+Y\cdot\nabla_{X}^{S}\varphi,$
    where $\nabla$ denotes the Levi-Civita connection on $(M,g)$. In case $n$ is even, the splitting $S^{+}\oplus S^{-}$ is $\nabla^{S}$-parallel.
The curvature $R^{S}$ of $\nabla^{S}$ is related to the Ricci curvature, $\operatorname{Ric}$, of $g$ via
$\sum_{k=1}^{n}\varepsilon_{k}e_{k}\cdot R^{S}(X,e_{k})\varphi=-\frac{1}{2}\operatorname{Ric}(X)\cdot\varphi+\frac{i}{2}(X\lrcorner\omega)\cdot\varphi,$ (2)
where $\{e_{1},\ldots,e_{n}\}$ is a pseudo-orthonormal frame with $\varepsilon_{k}=g(e_{k},e_{k})$ and $i\omega\in\Omega^{2}(M,i\mathbb{R})$ denotes the curvature 2-form coming from the connection 1-form $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$ [17, Page 65].
Remark 2.
The reason why we only consider time oriented manifolds, is to be able to define the above bundle metric. It can not be defined unless this assumption is included, see Remark 6.2.14 of [5].
3 Generalized Killing spinors
Let $(M,g)$ denote a time oriented semi-Riemannian $\text{spin}^{c}$ manifold of dimension $n$ and index $r$, with all the associated structure from the above section. For clarity, we repeat the definition from the introduction. A non-trivial spinor field $\varphi\in\Gamma(S)$ is said to be a generalized Killing spinor, if there is some $g$-symmetric endomorphism field $A$ of $TM$ such that
$\nabla^{S}_{X}\varphi=jA(X)\cdot\varphi\hskip 14.22636pt\forall X\in\Gamma(TM),$ (3)
where $j\in\{1,i\}$. We call $A$ the Killing endomorphism and if $j=1$ $(j=i)$ we say that $\varphi$ is a real (imaginary) generalized Killing spinor. In case $A=\mu\operatorname{Id}$ for some real valued function $\mu\in C^{\infty}(M)$, we say in addition that $\varphi$ is a special generalized Killing spinor with Killing function $j\mu$. Lastly, if $\mu$ is a real constant, we say that $\varphi$ is a Killing spinor with Killing number $j\mu$. If $\varphi$ is a generalized Killing spinor, we will simply say that it is special to mean that it is a special generalized Killing spinor.
Given a spinor field $\varphi\in\Gamma(S)$, we define the Dirac current $V_{\varphi}\in\Gamma(TM)$ as the vector field that is metrically equivalent to the $1$-form $\omega_{\varphi}:X\mapsto i^{r+1}\langle{\varphi,X\cdot\varphi}\rangle$. When $M$ has even dimension, we write $\varphi=\varphi^{+}+\varphi^{-}$ and use the notation $\overline{\varphi}:=\varphi^{+}-\varphi^{-}$ to define the one-form
$\overline{\omega}_{\varphi}:X\mapsto\langle{\varphi,X\cdot\overline{\varphi}}\rangle.$ (4)
Using the symmetry property $\langle{X\cdot\psi_{1},\psi_{2}}\rangle=(-1)^{r+1}\langle{\psi_{1},X\cdot\psi_{2}}\rangle$ mentioned in 2 of structures on the spinor bundle, we see that when the index $r$ is even, the above one-form is real valued, and when $r$ is odd, this one-form is imaginary valued. When $r$ is even, we define the modified Dirac current $\overline{V}_{\varphi}$ as the metrically equivalent vector field associated to $\overline{\omega}_{\varphi}$. When $r$ is odd, we define the modified Dirac current, $\overline{V}_{\varphi}$ as the metrically equivalent vector field to the $1$-form $i\overline{\omega}_{\varphi}$.
We say that a vector field is closed if its associated $1$-form is closed and we say that a vector field $V$ is conformal if it satisfies $L_{V}g=\psi g$, for some function $\psi\in C^{\infty}(M)$. Recall that a vector field $V$ is closed if and only if $g(\nabla_{X}V,Y)=g(\nabla_{Y}V,X)$, for all $X,Y\in TM$. Note that if $V$ is closed, then we have $(L_{V}g)(X,Y)=2g(\nabla_{X}V,Y)$, for all $X,Y\in TM$.
Lemma 3.
Let $\varphi\in\Gamma(S)$ be a generalized Killing spinor and suppose that either
  1. 1.
    the index $r$ is odd, and $j=1$, or
  2. 2.
    the index $r$ is even, and $j=i$, or
  3. 3.
    the index $r$ is odd, the dimension $n$ is even, and $j=i$, or
  4. 4.
    the index $r$ is even, the dimension $n$ is even, and $j=1$.
Then in Case 1 and Case 2, $V_{\varphi}$ is a closed vector field with
$\nabla_{X}V_{\varphi}=c\langle{\varphi,\varphi}\rangle A(X),$ (5)
where c=2c=-2 if $r\mod{4}\in\{0,3\}$ and $c=2$ otherwise. In Case 3 and Case 4, $\overline{V}_{\varphi}$ is a closed vector field with
$\nabla_{X}\overline{V}_{\varphi}=2\langle{\varphi,\overline{\varphi}}\rangle A(X).$ (6)
If $\varphi$ is furthermore a special generalized Killing spinor with $A=\mu\operatorname{Id}$, then in Case 1 and Case 2, $V_{\varphi}$ is conformal with $L_{V_{\varphi}}g=2c\mu\langle{\varphi,\varphi}\rangle g$. In Case 3 and Case 4, $\overline{V}_{\varphi}$ is conformal with $L_{\overline{V}_{\varphi}}g=4\mu\langle{\varphi,\overline{\varphi}}\rangle g$
Proof.
Given a generalized Killing spinor $\varphi\in\Gamma(S)$, we compute for vector fields $X$ and $Y$ that
$g(\nabla_{X}V_{\varphi},Y)$ $=X(g(V_{\varphi},Y))-g(V_{\varphi},\nabla_{X}Y)$
$=i^{r+1}(X(\langle{\varphi,Y\cdot\varphi}\rangle)-\langle{\varphi,\nabla_{X}Y\cdot\varphi}\rangle)$
$=i^{r+1}(\langle{\nabla_{X}^{S}\varphi,Y\cdot\varphi}\rangle+\langle{\varphi,\nabla_{X}^{S}(Y\cdot\varphi)}\rangle-\langle{\varphi,\nabla_{X}Y\cdot\varphi}\rangle)$
$=i^{r+1}(\langle{\nabla_{X}^{S}\varphi,Y\cdot\varphi}\rangle+\langle{\varphi,Y\cdot\nabla_{X}^{S}\varphi}\rangle)$
$=i^{r+1}(\langle{jA(X)\cdot\varphi,Y\cdot\varphi}\rangle+\langle{\varphi,Y\cdot jA(X)\cdot\varphi}\rangle).$ (7)
First, suppose we are in Case 1, i.e. the index $r$ is odd, and the generalized Killing spinor is real. Then by (3) we get that
$g(\nabla^{S}_{X}V_{\varphi},Y)$ $=i^{r+1}(\langle{A(X)\cdot\varphi,Y\cdot\varphi}\rangle+\langle{\varphi,Y\cdot A(X)\cdot\varphi}\rangle)$
$=i^{r+1}(\langle{\varphi,A(X)\cdot Y\cdot\varphi}\rangle+\langle{\varphi,Y\cdot A(X)\cdot\varphi}\rangle)$
$=-2i^{r+1}\langle{\varphi,\varphi}\rangle g(A(X),Y)$
$=c\langle{\varphi,\varphi}\rangle g(A(X),Y)$
which shows (5). Note that since $A$ is $g$-symmetric, we have also showed that $V_{\varphi}$ is closed. Using this, the conformal statement follows straightforwardly. Case 2 is obtained by a similar calculation. Indeed, we use (3) to get that
$g(\nabla_{X}V_{\varphi},Y)$ $=i^{r+1}(\langle{iA(X)\cdot\varphi,Y\cdot\varphi}\rangle+\langle{\varphi,Y\cdot iA(X)\cdot\varphi}\rangle)$
$=i^{r+2}(\langle{A(X)\cdot\varphi,Y\cdot\varphi}\rangle-\langle{\varphi,Y\cdot A(X)\cdot\varphi}\rangle)$
$=-i^{r+2}(\langle{\varphi,A(X)\cdot Y\cdot\varphi}\rangle+\langle{\varphi,Y\cdot A(X)\cdot\varphi}\rangle)$
$=2i^{r+2}\langle{\varphi,\varphi}\rangle g(A(X),Y)$
$=c\langle{\varphi,\varphi}\rangle g(A(X),Y),$
which completes Case 2. Suppose now that we are in Case 3. We note that $\nabla_{X}^{S}\overline{\varphi}=-A(X)\cdot\overline{\varphi}$ and use this to compute that
$g(\nabla_{X}\overline{V}_{\varphi},Y)$ $=X(g(\overline{V}_{\varphi},Y))-g(\overline{V}_{\varphi},\nabla_{X}Y)$
$=i(X(\langle{\varphi,Y\cdot\overline{\varphi}}\rangle)-\langle{\varphi,\nabla_{X}Y\cdot\overline{\varphi}}\rangle)$
$=i(\langle{\nabla^{S}_{X}\varphi,Y\cdot\overline{\varphi}}\rangle+\langle{\varphi,\nabla^{S}_{X}(Y\cdot\overline{\varphi})}\rangle-\langle{\varphi,\nabla_{X}Y\cdot\overline{\varphi}}\rangle)$
$=i(\langle{iA(X)\cdot\varphi,Y\cdot\overline{\varphi}}\rangle+\langle{\varphi,Y\cdot\nabla^{S}_{X}\overline{\varphi}}\rangle)$
$=-\langle{A(X)\cdot\varphi,Y\cdot\overline{\varphi}}\rangle-i\langle{\varphi,Y\cdot iA(X)\cdot\overline{\varphi}}\rangle$
$=-\langle{\varphi,A(X)\cdot Y\cdot\overline{\varphi}}\rangle-\langle{\varphi,Y\cdot A(X)\cdot\overline{\varphi}}\rangle)$
$=2\langle{\varphi,\overline{\varphi}}\rangle g(A(X),Y),$
which completes Case 3. Similarly, for Case 4, we compute that
$g(\nabla_{X}\overline{V}_{\varphi},Y)$ $=X(g(\overline{V}_{\varphi},Y))-g(\overline{V}_{\varphi},\nabla_{X}Y)$
$=X(\langle{\varphi,Y\cdot\overline{\varphi}}\rangle)-\langle{\varphi,\nabla_{X}Y\cdot\overline{\varphi}}\rangle$
$=\langle{\nabla^{S}_{X}\varphi,Y\cdot\overline{\varphi}}\rangle+\langle{\varphi,\nabla^{S}_{X}(Y\cdot\overline{\varphi})}\rangle-\langle{\varphi,\nabla_{X}Y\cdot\overline{\varphi}}\rangle$
$=\langle{A(X)\cdot\varphi,Y\cdot\overline{\varphi}}\rangle+\langle{\varphi,Y\cdot\nabla^{S}_{X}\overline{\varphi}}\rangle$
$=\langle{A(X)\cdot\varphi,Y\cdot\overline{\varphi}}\rangle-\langle{\varphi,Y\cdot A(X)\cdot\overline{\varphi}}\rangle$
$=-(\langle{\varphi,A(X)\cdot Y\cdot\overline{\varphi}}\rangle+\langle{\varphi,Y\cdot A(X)\cdot\overline{\varphi}}\rangle)$
$=2\langle{\varphi,\overline{\varphi}}\rangle g(A(X),Y),$
which completes the proof. ∎
Our goal is to utilize these closed and sometimes conformal vector fields to draw geometric conclusions of the underlying manifold.
4 The geometry of semi-Riemannian manifolds carrying closed and conformal vector fields
Closed and conformal vector fields on semi-Riemannian manifolds have been studied by Kühnel and Rademacher in [29] and [30]. In [29], they show the following. Note that they are considering gradient vector fields, which are locally the same as closed vector fields.
Lemma 4 (Kühnel & Rademacher, 1994, [29]).
Let $(M,g)$ be a semi-Riemannian manifold admitting a non-trivial conformal gradient field $V=\operatorname{grad}\psi$. Then the following holds.
  1. 1.
    In a neighborhood $U$ of any point $p\in M$ with $g(V,V)_{p}\neq 0$, we have that $(U,g|_{U})$ is isometric to a warped product of the form $(I\times F,\epsilon dr^{2}+(\psi^{\prime})^{2}g|_{F})$, where $\epsilon=\text{sign}(g(V,V))$ and the submanifolds $\{t\}\times F$ corresponds to the level sets of $\psi$. Furthermore, the integral curves of $\frac{V}{|V|}$ are geodesics and the function $\psi:I\to\mathbb{R}$ satisfies $\psi^{\prime\prime}=\epsilon\frac{\Delta\psi}{n}$, where $\Delta$ denotes the Laplacian.
  2. 2.
    The zeroes of $V$ are isolated. For each such zero $p\in M$, there exists a normal neighborhood $U\subset T_{p}M$, for which there are coordinates on $U\setminus C_{p}$, with $C_{p}$ denoting the light cone, where the metric takes the form $g_{(r,x)}=\epsilon dr^{2}+(\frac{\psi_{\epsilon}^{\prime}(r)}{\psi_{\epsilon}^{\prime\prime}(0)})^{2}\tilde{g}_{x}$, with $\epsilon=\text{sign}\hskip 2.84526ptg(V,V)$, and $\tilde{g}$ denotes the standard metric on $\{x\in\mathbb{R}^{n}\mid\langle{x,x}\rangle_{r,n-r}=\epsilon\}$. In these coordinates, we have that $\psi_{\epsilon}(r)=\psi(r,x)$. Furthermore, these coordinates extend to a conformally flat metric on all of $U$.
Stronger results in the Riemannian setting have been proved by Rademacher in [37]. Our goal in this section is to strengthen the result 1 of Lemma 4 above. We intend to do so by expanding on the ideas and techniques developed by Olea and Gutiérrez in [25] and [26]. First, we proceed without assuming completeness.
4.1 Non-local splittings without completeness assumptions
Throughout, we assume that $(M,g)$ is a connected semi-Riemannian manifold of dimension at least $2$.
Let $V$ be a closed vector field with nowhere vanishing norm $|V|:=\sqrt{|g(V,V)|}$ and put $E:=\frac{V}{|V|^{2}}$. Let $\Phi:D\to M$, with $D\subset\mathbb{R}\times M$, be the unique maximal flow of $E$. Recall that $E$ is said to be a complete vector field if $D=\mathbb{R}\times M$. For each $t\in\mathbb{R}$, we get a map $\Phi_{t}:M_{t}\to M_{-t}$, where $M_{t}:=\{p\in M\mid(t,p)\in D\}$. The map $\Phi_{t}$ is a diffeomorphism with inverse $\Phi_{-t}$. In case $V$ is also conformal, we will also consider the maximal flow $\tilde{\Phi}$ of the vector field $\tilde{E}:=\frac{V}{|V|}$. We let $\tilde{D}\subset\mathbb{R}\times M$ denote the domain of $\tilde{\Phi}$.
One reason why we are interested in closed vector fields $V$ is that the distribution $V^{\perp}\subset TM$ is integrable and thus induces a foliation $\mathcal{F}=\{L_{p}\}_{p\in M}$ of codimension $1$ submanifolds $L_{p}$.
In [25] and [26] it is assumed in addition that $\tilde{E}$ is complete. This yields global results, which we will come back to later. To clarify matters in the non-complete case, we need to introduce the following definition. We say that a connected open subset $L_{p}^{\prime}\subset L_{p}$ is a uniform subleaf with respect to the flow $\Phi$ ($\tilde{\Phi}$), if there exists some non-empty open interval $I$, containing 0, such that $I\times L_{p}^{\prime}\subset D$ ($I\times L_{p}^{\prime}\subset\tilde{D}$). We call the pair $(L_{p}^{\prime},I)$ a uniform pair for $\Phi$ ($\tilde{\Phi}$). Given a uniform pair $(I,L_{p}^{\prime})$ for $\Phi$ ($\tilde{\Phi}$), we write $\Phi^{L_{p}^{\prime}}$ ($\tilde{\Phi}^{L_{p}^{\prime}}$) to denote the restriction of $\Phi$ to $I\times L_{p}^{\prime}$ ($\tilde{\Phi}$ to $I\times L_{p}$). This notation will be unambiguous since the interval $I$ will be clear from context. Note that the point $p$ might not be a point of $L_{p}^{\prime}$. However, for any $q\in L_{p}$, we have that $L_{q}=L_{p}$, and therefore we will always assume that $p$ is chosen so that $p\in L_{p}^{\prime}$.
We will sometimes use the following fact, which follows straightforwardly from the definition of a flow. Given a uniform pair $(I,L_{p}^{\prime})$ for $\Phi$, the map $\Phi^{L_{p}^{\prime}}$ is injective if and only if each integral curve for $E$ starting at a point on $L_{p}^{\prime}$ never returns to $L_{p}^{\prime}$.
Assume that $V$ is also conformal with $L_{V}g=\psi g$. Then it follows from the definitions that $\psi=\frac{2\operatorname{div}V}{n}$ and $\nabla_{X}V=\frac{\operatorname{div}V}{n}X$, for all $X\in TM$. By the last property, $|V|$ is constant along the leaves $L_{p}$. We collect some straightforward formulas which will be used later. We compute for any $X,Y\perp V$ that
$(L_{\tilde{E}}g)(X,Y)=\frac{(L_{V}g)(X,Y)}{|V|}=\frac{2\operatorname{div}V}{n|V|}.$ (8)
Using the formula $\nabla_{X}V=\frac{\operatorname{div}V}{n}X$ from above, we obtain that
$\tilde{E}(|V|)=\frac{\operatorname{div}V}{n}.$ (9)
We say that $\Phi$ ($\tilde{\Phi}$) preserves the foliation $\mathcal{F}$, if for every uniform pair $(I,L_{p}^{\prime})$ for $\Phi$ ($\tilde{\Phi}$), we have that $\Phi_{t}(L_{p}^{\prime})\subset L_{\Phi_{t}(p)}$ ($\tilde{\Phi}_{t}(L_{p}^{\prime})\subset L_{\tilde{\Phi}_{t}(p)}$).
Proposition 5.
Let $V$ be a closed vector field with nowhere vanishing norm and use the notation from above. Then $\Phi$ preserves the foliation $\mathcal{F}$. If $V$ is also conformal, then $\tilde{\Phi}$ also preserves the foliation $\mathcal{F}$.
Proof.
Fix some uniform pair $(I,L_{p}^{\prime})$ for $\Phi$. First, assume that $\Phi|_{I\times L_{p}^{\prime}}$ is injective. Take some point $q\in L_{p}^{\prime}$, with $q\neq p$. Since $L_{p}^{\prime}$ is connected, there exists a path with embedded image $\ell\cong(0,1)$ in $L_{p}^{\prime}$, containing both the point $p$ and the point $q$. Take a tubular neighborhood $\mathcal{U}$ of $\ell$ in $L_{p}^{\prime}$. Since $V$ is nowhere tangent to $\mathcal{U}$ and $\Phi|_{I\times\hskip 0.85358pt\mathcal{U}}$ is injective, we have that $\Phi|_{I\times\hskip 0.85358pt\mathcal{U}}$ is a diffeomorphism onto an open submanifold $N$ of $M$ [32, Theorem 9.20]. Since $N$ has trivial de Rahm cohomology, we can write $V|_{N}=\operatorname{grad}h$, for some function $h:N\to\mathbb{R}$ and note that the intersection of each leaf $L_{m}\in\mathcal{F}$ with $N$ is either empty or a level set of $h$. Hence, if we show that $h(\Phi_{t}(p))=h(\Phi_{t}(q))$, for all $t\in I$, then $\Phi_{t}(q)\in L_{\Phi_{t}(p)}$ and we could conclude the current case. By assumption, it is true for $t=0$. Therefore, it suffices to show that $\frac{d}{dt}h(\Phi_{t}(p))=\frac{d}{dt}h(\Phi_{t}(q))$ for all $t\in I$. We compute that
$\frac{d}{dt}h(\Phi_{t}(p))$ $=dh(E_{\Phi_{t}(p)})=g(V,\frac{V}{|V|^{2}})_{\Phi_{t}(p)}=\operatorname{sign}{g(V,V)}.$ (10)
With the same computation, we see that $\frac{d}{dt}h(\Phi_{t}(q))=\operatorname{sign}{g(V,V)}$ and we can conclude the case when $\Phi|_{I\times L_{p}^{\prime}}$ is injective. Now, suppose instead that $\Phi|_{I\times L_{p}^{\prime}}$ is not injective. Let $T\in\mathbb{R}$ be the largest positive real number such that $\Phi|_{(I\cap(-T,T))\times L_{p}^{\prime}}$ is injective. Hence, there exists an integral curve for $E$, starting at a point of $L_{p}^{\prime}$, which returns to $L_{p}^{\prime}$ either at $T$ or at T-T, and we assume without loss of generality that it returns at $T$. There exists some $\varepsilon>0$ such that $\Phi|_{(-\varepsilon,\varepsilon)\times\Phi_{T}(L_{p}^{\prime})}$ is injective, and so by the first part of the proof it follows that every integral curve for $E$ starting at a point of $L_{p}^{\prime}$ returns to $L_{p}$ at $T$. By induction, each integral curve starting at a point of $L_{p}^{\prime}$ returns to $L_{p}$ exactly at $t\in\{kT\}_{k\in\mathbb{Z}}\cap I$. Hence $\Phi|_{(((k-1)T,(k+1)T)\hskip 1.42262pt\cap\hskip 1.42262ptI)\times L_{p}^{\prime}}$ is injective for all $k$. It follows that $\Phi|_{((-T,T)\hskip 1.42262pt\cap\hskip 1.42262pt(I-kT))\times\Phi_{kT}(L_{p}^{\prime})}$ is injective, where we have used the notation (a,b)kT:=(akT,bkT)(a,b)-kT:=(a-kT,b-kT). For all $k\in\mathbb{Z}$, we apply the first part of the proof to $\Phi|_{((-T,T)\hskip 1.42262pt\cap\hskip 1.42262pt(I-kT))\times\Phi_{kT}(L_{p}^{\prime})}$ and conclude that for all $t\in(-T,T)\cap(I-kT)$, we have that $\Phi_{t+kT}(L_{p}^{\prime})=\Phi_{t}(\Phi_{kT}(L_{p}^{\prime}))\subset L_{\Phi_{t}(\Phi_{kT}(p))}=L_{\Phi_{t+kT}(p)}$, and it follows that $\Phi$ preserves the foliation. Assume that $V$ is also conformal. Then $|V|$ is constant along the leaves of the foliation $\mathcal{F}$. We can, except for the argument that $\frac{d}{dt}h(\Phi_{t}(p))=\frac{d}{dt}h(\Phi_{t}(q))$, use the same proof as above to show that $\tilde{\Phi}$ preserves the foliation $\mathcal{F}$. To show that $\frac{d}{dt}h(\Phi_{t}(p))=\frac{d}{dt}h(\Phi_{t}(q))$, we use that $V$ is constant along the leaves to compute that $\frac{d}{dt}h(\Phi_{t}(p))=(\operatorname{sign}{g(V,V)})|V|_{p}=(\operatorname{sign}{g(V,V)})|V|_{q}=\frac{d}{dt}h(\Phi_{t}(q))$ and thus we may complete the proof. ∎
Proposition 6.
Let $V$ be a closed vector field on $(M,g)$ with nowhere vanishing norm, and let $\Phi$ and $\tilde{\Phi}$ be given as above. Then, for each uniform pair $(L_{p}^{\prime},I)$ for $\Phi$, the map
$\Phi^{L_{p}^{\prime}}:(I\times L_{p}^{\prime},\frac{dt^{2}}{g(V,V)}+h_{t})\to(M,g),$
where $h_{t}=\Phi_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$, is a local isometry. Furthermore, if $V$ is also conformal, then for each uniform pair $(L_{p}^{\prime},I)$ for $\tilde{\Phi}$, the map
$\tilde{\Phi}^{L_{p}^{\prime}}:(I\times L_{p}^{\prime},\epsilon\hskip 0.73975ptdt^{2}+f^{2}g|_{L_{p}^{\prime}})\to(M,g),$
[…]
Proof.
By construction, $\Phi^{L_{p}^{\prime}}:I\times L_{p}^{\prime}\to M$ is a local diffeomorphism relating $\frac{\partial}{\partial t}$ with $\frac{V}{|V|^{2}}$. By Proposition 5, $\Phi$ preserves the foliation induced by $V^{\perp}$ and therefore the pullback metric $(\Phi^{L_{p}^{\prime}})^{*}(g)$ on $I\times L_{p}^{\prime}$ can be written as $\frac{dt^{2}}{g(V,V)}+h_{t}$, where $h_{t}=\Phi_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$. This proves the first part. Assume now that $V$ is also conformal. By an argument similar to the above, we get that the pullback metric $(\tilde{\Phi}^{L_{p}^{\prime}})^{*}(g)$ is of the form $\epsilon\hskip 0.73975ptdt^{2}+m_{t}$, where $m_{t}=\tilde{\Phi}_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$. It remains to determine $m_{t}$. Fix $v,w\in T_{x}(L_{p}^{\prime})$. We intend to show that $m_{t}(v,w)$ satisfies the differential equation $\frac{d}{dt}m_{t}(v,w)=(f^{2})^{\prime}m_{t}(v,w)$. Using (8), we compute that
$\frac{d}{dt}(m_{t}(v,w))$ $=\frac{d}{dt}((\tilde{\Phi}_{t}^{*}g)(v,w))$
$=(L_{\tilde{E}}g)((\tilde{\Phi}_{t})_{*}v,(\tilde{\Phi}_{t})_{*}w)$
$=\frac{2\operatorname{div}V}{n|V|}(\tilde{\Phi}_{t}(x))g((\tilde{\Phi}_{t})_{*}v,(\tilde{\Phi}_{t})_{*}w),$
$=\frac{2\operatorname{div}V}{n|V|}(\tilde{\Phi}_{t}(x))(m_{t}(v,w)),$
and hence the abovementioned differential equation is satisfied. Showing that $\frac{2\operatorname{div}V}{n|V|}(\tilde{\Phi}_{t}(x))$ is only a function of $t$, completes the proof. As stated above, $|V|$ is constant along the leaves. Since $\tilde{\Phi}^{L_{p}^{\prime}}$ is a local diffeomorphism, we have that around any point $p\in M$ there is a local frame of the form $\{\tilde{E},\partial_{1},\ldots,\partial_{n-1}\}$, where $\partial_{k}\perp V$ for all $k\in\{1,\ldots,n-1\}$. We use (9) and that $|V|$ is constant along the leaves, to compute that
$\partial_{k}\operatorname{div}V=n\partial_{k}\tilde{E}(|V|)=n\tilde{E}\partial_{k}(|V|)=0.$
Thus the function $\operatorname{div}V$ is constant along the leaves and consequently, so is $\frac{2\operatorname{div}V}{n|V|}(\tilde{\Phi}_{t}(x))$. Hence we may unambiguously write $\frac{2\operatorname{div}V}{n|V|}(\tilde{\Phi}_{t}(x))=\frac{2\operatorname{div}V}{n|V|}(t)$ and conclude the proof. ∎
When $V$ is both closed and conformal, one can give conditions for when the above local isometry is an isometry onto its image. This is the content of the following proposition, which is a generalization of a result due to Gutierrez and Olea in [25] and [26].
Proposition 7.
Let $V$ be a closed and conformal vector field on $(M,g)$ with nowhere vanishing norm, and let $\tilde{\Phi}$ be given as above. Further, let $(L_{p}^{\prime},I)$ be a uniform pair for $\tilde{\Phi}$ such that $\operatorname{div}V_{p}\neq 0$, and write $N:=\tilde{\Phi}^{L_{p}^{\prime}}(I\times L_{p}^{\prime})$. If either
  1. 1.
    $\operatorname{div}(V)|_{N}$ is non-positive or non-negative, or
  2. 2.
    $\operatorname{Ric}(V,V)|_{N}$ is non-positive or non-negative,
then the map $\tilde{\Phi}^{L_{p}^{\prime}}$ is injective.
Proof.
The map $\tilde{\Phi}^{L_{p}^{\prime}}$ is injective if and only if each integral curve of $\tilde{E}$, starting at some point of $L_{p}^{\prime}$, with image contained in $N$, intersects $L_{p}^{\prime}$ only at $t=0$. Fix an integral curve $\gamma$ of $\tilde{E}$ with $\gamma(0)=q\in L_{p}^{\prime}$ and image contained in $N$. Since $|V|$ is constant along the leafs of $\mathcal{F}$, the function $|V|_{\gamma(t)}$ only depends on which leaf $\gamma(t)$ is in. Also, by (9), the condition that $\operatorname{div}V_{p}\neq 0$ means that $\tilde{E}(|V|)_{q}\neq 0$. Hence the function $|V|_{\gamma(t)}$ is not constant.
First, suppose that $\operatorname{div}(V)|_{N}\geq 0$. By (9), we have that $\frac{d}{dt}|V|_{\gamma(t)}\geq 0$ and therefore $|V|_{\gamma(t)}$ is a non-constant monotone function. Hence, $\gamma(t)$ can never return to $L_{p}^{\prime}$ for any $t\neq 0$. The proof is completely analogous when $\operatorname{div}(V)|_{N}\leq 0$.
Secondly, using (9), we get that $\nabla_{X}V=\tilde{E}(|V|)X$, so we may compute that
$R(X,V)V=\nabla_{X}\nabla_{V}V-\nabla_{V}\nabla_{X}V-\nabla_{[X,V]}V=X(\tilde{E}(|V|))V-V(\tilde{E}(|V|))X,$
and therefore, $\operatorname{Ric}(V,V)=-(n-1)V(\tilde{E}(|V|))$. Assume that $\operatorname{Ric}(V,V)|_{N}$ is non-negative and suppose for a contradiction that there is some point $0\neq t_{0}\in I$ such that $\gamma(t_{0})\in L_{p}^{\prime}$. By our computation of $\operatorname{Ric}(V,V)$, the function $\frac{\partial^{2}}{\partial t^{2}}|V|_{\gamma(t)}$ is non-positive. By Proposition 5, the flow $\tilde{\Phi}$ preserves the foliation, and therefore we have that $|V|_{\gamma(t)}=|V|_{\gamma(t_{0}+t)}$ for all $t$ for which both sides are well defined. This means that we can extend the non-constant smooth function $|V|_{\gamma(\cdot)}$ to a non-constant smooth and periodic function $\mathbb{R}\to\mathbb{R}$ with period $t_{0}$, whose second derivative is non-positive. This is a contradiction since smooth non-constant periodic functions must have both positive and negative second derivatives. The same proof applies if we assume that $\operatorname{Ric}(V,V)|_{N}$ is non-positive. ∎
4.2 Global splittings under completeness assumptions
We proceed with the notation from the previous section. Given a flow $\Psi$ and an interval $(t_{1},t_{2})\subset\mathbb{R}$, we will sometimes write $\Psi_{(t_{1},t_{2})}(V)$, to denote $\Psi((t_{1},t_{2})\times V)$.
Proposition 8.
Let $\Psi$ be the flow of a nowhere vanishing complete vector field $W$ on a connected smooth manifold $P$, preserving a codimension $1$ foliation $\mathcal{Q}=\{Q_{p}\}_{p\in P}$. Then, for any leaf $Q_{p}\in\mathcal{Q}$, the map $\Psi|_{\mathbb{R}\times Q_{p}}:\mathbb{R}\times Q_{p}\to P$ is a normal covering map.
Proof.
Fix some $Q_{p}\in\mathcal{Q}$. First, we show that $\Psi|_{\mathbb{R}\times Q_{p}}:\mathbb{R}\times Q_{p}\to P$ is surjective. We will show that the image $\Psi(\mathbb{R}\times Q_{p})=:N$ is both open and closed. Since the vector field $W$ is nowhere vanishing, we have that $\Psi|_{\mathbb{R}\times Q_{p}}$ is a local diffeomorphism and thus, $N$ is open. If $N^{c}$ is empty, then we are done. Suppose $N^{c}$ is not empty and take some $x\in N^{c}$. It follows that $\Psi(\mathbb{R}\times Q_{x})\subset N^{c}$. Indeed, if there where points $t_{1},t_{2}\in\mathbb{R}$, $y_{1}\in Q_{x}$, $y_{2}\in Q_{p}$ such that $\Psi(t_{1},y_{1})=\Psi(t_{2},y_{2})$, then we can combine the integral curves $\Psi(-,y_{1})$ and $\Psi(-,y_{2})$ to obtain an integral curve from $Q_{p}$ to $Q_{x}$, which would mean that $x\in N$. Hence, $x$ is contained in the open subset $\Psi(\mathbb{R}\times Q_{x})$, which is a subset of $N^{c}$. This holds for all points in $N^{c}$ and thus $N^{c}$ is open. Hence $\Psi|_{\mathbb{R}\times Q_{p}}:\mathbb{R}\times Q_{p}\to P$ is surjective. Now, we will show that every $x\in P$ has an evenly covered neighborhood. If $\Psi|_{\mathbb{R}\times Q_{p}}$ is injective, then we are done because $\Psi|_{\mathbb{R}\times Q_{p}}$ would be a diffeomorphism. Suppose therefore that $\Psi|_{\mathbb{R}\times Q_{p}}$ is not injective. Take some integral curve $\gamma$ of $W$ starting at a point of $Q_{p}$. Let $T\in\mathbb{R}$ be the smallest positive real number such that $\gamma(T)\in Q_{p}$. Since the flow $\Psi$ preserves the foliation $\mathcal{Q}$, we have that each integral curve of $W$ starting at a point on $Q_{p}$ will return to the leaf $Q_{p}$ at exactly the times $\{kT\mid k\in\mathbb{Z}\}$. Fix some point $x\in P$. Since $W$ is nowhere vanishing, there exists some $\varepsilon>0$ such that $\Psi|_{(-\varepsilon,\varepsilon)\times Q_{x}}:(-\varepsilon,\varepsilon)\times Q_{x}\to\Psi((-\varepsilon,\varepsilon)\times Q_{x})=:U$ is a diffeomorphism. Hence $U$ is open. Let $t_{x}\in\mathbb{R}$ be the smallest positive real number such that $\Psi_{t_{x}}(Q_{p})=Q_{x}$. We note that
$\Psi((t_{x}+kT-\varepsilon,t_{x}+kT+\varepsilon)\times Q_{p})$ $=\Psi_{(t_{x}+kT-\varepsilon,t_{x}+kT+\varepsilon)}(Q_{p})$
$=\Psi_{(-\varepsilon,\varepsilon)}\circ\Psi_{t_{x}}\circ\Psi_{kT}(Q_{p})$
$=\Psi_{(-\varepsilon,\varepsilon)}\circ\Psi_{t_{x}}(Q_{p})$
$=\Psi_{(-\varepsilon,\varepsilon)}(Q_{x})=U,$
and by the minimality assumptions on $T$ and $t_{x}$, we conclude that
$(\Psi|_{\mathbb{R}\times Q_{p}})^{-1}(U)=\bigcup_{k\in\mathbb{Z}}((t_{x}+kT-\varepsilon,t_{x}+kT+\varepsilon)\times Q_{p}).$
Hence, $\Psi|_{\mathbb{R}\times Q_{p}}$ is a covering map. We show that the group of deck transformations of the covering act transitively on the fibers. Take any $(t_{0},p_{0}),(t_{1},p_{1})\in\mathbb{R}\times Q_{p}$ in the fiber over some $y\in P$. We claim that the diffeomorphism
$f:\mathbb{R}\times Q_{p}$ $\to\mathbb{R}\times Q_{p}$
$(t,m)$ $\mapsto(t+(t_{0}-t_{1}),\Psi((t_{1}-t_{0}),m))$
is a deck transformation mapping $(t_{1},p_{1})$ to $(t_{0},p_{0})$. It is a deck transformation since
$\Psi(f(t,m))$ $=\Psi(t+(t_{0}-t_{1}),\Psi((t_{1}-t_{0}),m))$
$=\Psi_{t}\circ\Psi((t_{0}-t_{1}),\Psi((t_{1}-t_{0}),m))$
$=\Psi_{t}\circ\Psi(0,m)=\Psi(t,m),$
and it maps $(t_{1},p_{1})$ to $(t_{0},p_{0})$ since
$f(t_{1},p_{1})$ $=(t_{0},\Psi((t_{1}-t_{0}),p_{1}))=(t_{0},\Psi_{-t_{0}}\circ\Psi(t_{1},p_{1}))=(t_{0},\Psi_{-t_{0}}(y))=(t_{0},p_{0}).$
Hence we conclude that $\Psi|_{\mathbb{R}\times Q_{p}}$ is a normal covering map. ∎
With this, we improve Proposition 6 under the additional assumption of $E$ or $\tilde{E}$ being complete.
Proposition 9.
Let $V$ be a closed vector field on $(M,g)$ with nowhere vanishing norm, and let $E$, $\Phi$, $\tilde{E}$, and $\tilde{\Phi}$ be given as in the previous section. If $E$ is complete, then for any leaf $L_{p}$ of $\mathcal{F}$, we have that
$\Phi^{L_{p}}:(\mathbb{R}\times L_{p},\frac{dt^{2}}{g(V,V)}+h_{t})\to(M,g),$
where $h_{t}=\Phi_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$, is a normal, semi-Riemannian covering. Assume in addition that $V$ is also conformal. If $\tilde{E}$ is complete, then for any leaf $L_{p}$ of $\mathcal{F}$, we have that
$\tilde{\Phi}^{L_{p}}:(\mathbb{R}\times L_{p},\epsilon\hskip 0.73975ptdt^{2}+f^{2}g|_{L_{p}})\to(M,g),$
[…]
Proof.
The statement follows by Proposition 8 and Proposition 6. ∎
By applying Proposition 7 we obtain the following corollary.
Corollary 10.
Let $V$ be a closed and conformal vector field on $(M,g)$ with nowhere vanishing norm. Suppose $\tilde{E}$ is complete, and $\operatorname{div}V$ is not constantly zero. If either $\operatorname{div}V$ or $\operatorname{Ric}(V,V)$ is non-positive or non-negative, then the covering $\tilde{\Phi}^{L_{p}}$ from Proposition 9 is an isometry for all leaves $L_{p}$.
Proof.
Let $p$ be such that $\operatorname{div}V_{p}\neq 0$. By Proposition 7, we have that $\tilde{\Phi}^{L_{p}}$ is an isometry. By Proposition 5, $\tilde{\Phi}$ preserves the foliation, and therefore the same is true for any other leaf $L_{q}\in\mathcal{F}$. ∎
5 Returning to generalized Killing spinors
We will apply the results of the previous section to semi-Riemannian $\text{spin}^{c}$ manifolds carrying generalized Killing spinors. In the Riemannian case, we have a natural distinction of type I and type II imaginary generalized Killing spinors, and this allows us to further develop the classification. Given a type I imaginary generalized Killing spinor on a Riemannian $\text{spin}^{c}$ manifold, we will find an explicit parallel spinor on each leaf of the foliation given by the orthogonal complement of the Dirac current. One could make ad hoc conditions for when similar results hold in the semi-Riemannian setting, but we do not conform to this. Nevertheless, due to a result by Bohle [11], we can find Killing and sometimes parallel spinors on the leaves when we restrict our attention to Killing spinors on time oriented semi-Riemannian spin manifolds.
5.1 The semi-Riemannian $\text{spin}^{c}$ setting
Let $(M,g)$ be a connected time oriented semi-Riemannian $\text{spin}^{c}$ manifold of index $r$ and dimension $n$. Let $\varphi$ be a generalized Killing spinor as in either of the cases of Lemma 3. To simplify notation, we shall throughout the remainder of this section write $V$ to denote either $V_{\varphi}$, corresponding to Case 1 and Case 2 of Lemma 3, or $\overline{V}_{\varphi}$, corresponding to Case 3 and Case 4 of Lemma 3. Hence, by Lemma 3, $V$ is a closed vector field and if $\varphi$ is a special generalized Killing spinor, then $V$ is also conformal. Let $\mathcal{F}=\{L_{p}\}_{p\in M}$ be the foliation induced by $V^{\perp}$. Further, we write $\beta$ to denote either $\langle{\varphi,\varphi}\rangle$ or $\langle{\varphi,\overline{\varphi}}\rangle$, with the analog case distinction as for $V$. Since $\nabla_{X}V=c\beta A(X)$, we get that $\operatorname{div}V=c\beta\operatorname{tr}(A)$, where in Case 1 and Case 2, $c\in\{\pm 2\}$ corresponds to the constant $c$ given in Lemma 3, and in Case 3 and Case 4, $c=2$. If $|V|$ is nowhere vanishing, then we put $E:=\frac{V}{|V|^{2}}$ and let $\Phi$ be the unique maximal flow of $E$. Similarly, let $\tilde{\Phi}$ be the unique maximal flow of $\tilde{E}:=\frac{V}{|V|}$.
The following facts are straightforward computations. In Case 1 and Case 4, the function $Q_{+}=\beta^{2}+g(V,V)$ is locally constant. In Case 2 and Case 3, the function $Q_{-}=\beta^{2}-g(V,V)$ is locally constant. Hence, in the corresponding cases, we have that $|V|=\sqrt{|Q_{\pm}-\beta^{2}|}$. If $|V|$ is nowhere vanishing, we shall write $\tilde{\beta}:=c\frac{\beta}{|V|}=c\frac{\beta}{\sqrt{|Q_{\pm}-\beta^{2}|}}$ and $\epsilon=\text{sign}(g(V,V))$, to simplify the formulas below.
Given a hypersurface and a unit normal $\xi$, we write $W^{\xi}$ to denote the Weingarten map $X\mapsto-(\nabla_{X}\xi)^{\perp}$.
Proposition 11.
Let $(M,g)$ be a time oriented semi-Riemannian $\text{spin}^{c}$ manifold of index $r$ and dimension $n$ carrying a generalized Killing spinor $\varphi$, with $|V|$ nowhere vanishing, such that any of the cases of Lemma 3 holds. For any leaf $L_{p}$, the Weingarten map $W^{\tilde{E}}$ for $L_{p}$ can be written as
[…]
For any uniform pair $(L_{p}^{\prime},I)$ for $\Phi$, the map $\Phi^{L_{p}^{\prime}}:(I\times L_{p}^{\prime},\frac{dt^{2}}{g(V,V)}+h_{t})\to(M,g)$, where $h_{t}=\Phi_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$, is a local isometry. If $\varphi$ is special with $A=\mu\operatorname{Id}$, then for any uniform pair $(L_{p}^{\prime},I)$ for $\tilde{\Phi}$ we get the local isometry $\tilde{\Phi}^{L_{p}^{\prime}}:(I\times L_{p}^{\prime},\epsilon\hskip 0.73975ptdt^{2}+f^{2}g|_{L_{p}^{\prime}})\to(M,g)$, where
$f(t)=\exp{\int_{0}^{t}\mu\tilde{\beta}ds}.$ (12)
Furthermore, if the subleaf $L_{p}^{\prime}$ is such that $(\mu\beta)_{p}\neq 0$ and the interval $I$ is such that either the function $\mu\beta$ or the function
  • $4(n-1)j^{2}\mu^{2}g(V,V)+2\overline{j}i^{r+1}(nV(\mu)\langle{\varphi,\varphi}\rangle+\langle{\varphi,\operatorname{grad}(\mu)\cdot V\cdot\varphi}\rangle)$ in Case 1 and Case 2
  • $4(n-1)j^{2}\mu^{2}g(V,V)-2nV(\mu)\langle{\varphi,\overline{\varphi}}\rangle-2\langle{\operatorname{grad}(\mu)\cdot V\cdot\varphi,\overline{\varphi}}\rangle$     in Case 3 and Case 4
is non-negative or non-positive on $\tilde{\Phi}(I\times L_{p}^{\prime})$, then the map $\tilde{\Phi}^{L_{p}^{\prime}}$ is injective.
Proof.
We compute that
$\nabla_{X}\tilde{E}$ $=X(|g(V,V)|^{-\frac{1}{2}})V+|V|^{-1}\nabla_{X}V$
$=-\frac{1}{2}|g(V,V)|^{-\frac{3}{2}}2\epsilon g(\nabla_{X}V,V)V+|V|^{-1}c\beta A(X)$
$=-|g(V,V)|^{-\frac{3}{2}}c\beta\epsilon g(A(X),V)V+|V|^{-1}c\beta A(X)$
[…]
and hence we obtain the formula for the Weingarten map. The statements about the local isometries follows by Lemma 3 and Proposition 6. The injectivity statement is a translation of the conditions of Proposition 7 into this setting. By the above, we get that $\operatorname{div}V=cn\mu\beta$. To compute $\operatorname{Ric}(V,V)$, we first compute that
$R^{S}(X,e_{k})\varphi$ $=\nabla^{S}_{X}\nabla^{S}_{e_{k}}\varphi-\nabla^{S}_{e_{k}}\nabla^{S}_{X}\varphi-\nabla^{S}_{[X,e_{k}]}\varphi$
$=j(\nabla^{S}_{X}\mu e_{k}\cdot\varphi-\nabla^{S}_{e_{k}}\mu X\cdot\varphi-\mu[X,e_{k}]\cdot\varphi)$
$=j(X(\mu)e_{k}\cdot\varphi+\mu e_{k}\cdot\nabla^{S}_{X}\varphi-e_{k}(\mu)X\cdot\varphi-\mu X\cdot\nabla^{S}_{e_{k}}\varphi)$
$=j(X(\mu)e_{k}\cdot\varphi+j\mu^{2}e_{k}\cdot X\cdot\varphi-e_{k}(\mu)X\cdot\varphi-j\mu^{2}X\cdot e_{k}\cdot\varphi)$
$=j(X(\mu)e_{k}-e_{k}(\mu)X)\cdot\varphi+j^{2}\mu^{2}(e_{k}\cdot X\cdot\varphi-X\cdot e_{k}\cdot\varphi),$
and hence the left hand side of (2) becomes
$LHS$ $=\sum_{k=1}^{n}\varepsilon_{k}e_{k}\cdot(j(X(\mu)e_{k}-e_{k}(\mu)X)\cdot\varphi+j^{2}\mu^{2}(e_{k}\cdot X\cdot\varphi-X\cdot e_{k}\cdot\varphi))$
$=-jnX(\mu)\varphi-j\operatorname{grad}(\mu)\cdot X\cdot\varphi+2j^{2}\mu^{2}\sum_{k=1}^{n}\varepsilon_{k}e_{k}\cdot(e_{k}\cdot X\cdot\varphi+g(X,e_{k})\varphi)$
$=-jnX(\mu)\varphi-j\operatorname{grad}(\mu)\cdot X\cdot\varphi-2(n-1)j^{2}\mu^{2}X\cdot\varphi,$
and thus we can write the identity (2) as
$\operatorname{Ric}(X)\cdot\varphi=2jnX(\mu)\varphi+2j\operatorname{grad}(\mu)\cdot X\cdot\varphi+4(n-1)j^{2}\mu^{2}X\cdot\varphi+i(X\lrcorner\omega)\cdot\varphi.$
With this, we compute in Case 1 and Case 2 that
$\operatorname{Ric}(V,V)$ $=g(\operatorname{Ric}(V),V)$
$=i^{r+1}\langle{\varphi,\operatorname{Ric}(V)\cdot\varphi}\rangle$
$=i^{r+1}\langle{\varphi,2jnV(\mu)\varphi+2j\operatorname{grad}(\mu)\cdot V\cdot\varphi+4(n-1)j^{2}\mu^{2}V\cdot\varphi+i(V\lrcorner\omega)\cdot\varphi}\rangle$
$=4(n-1)j^{2}\mu^{2}g(V,V)+2\overline{j}i^{r+1}(nV(\mu)\langle{\varphi,\varphi}\rangle+\langle{\varphi,\operatorname{grad}(\mu)\cdot V\cdot\varphi}\rangle),$
where in the last step we used that $\omega$ is a 2-form. Similarly, for Case 3 we compute that
$\operatorname{Ric}(V,V)$ $=g(\operatorname{Ric}(V),V)=i\langle{\varphi,\operatorname{Ric}(V)\cdot\overline{\varphi}}\rangle=i\langle{\operatorname{Ric}(V)\cdot\varphi,\overline{\varphi}}\rangle$
$=i\langle{2inV(\mu)\varphi+2i\operatorname{grad}(\mu)\cdot V\cdot\varphi-4(n-1)\mu^{2}V\cdot\varphi+i(V\lrcorner\omega)\cdot\varphi,\overline{\varphi}}\rangle$
$=-4(n-1)\mu^{2}g(V,V)-2\langle{nV(\mu)\varphi+\operatorname{grad}(\mu)\cdot V\cdot\varphi,\overline{\varphi}}\rangle$
$=-4(n-1)\mu^{2}g(V,V)-2\langle{nV(\mu)\varphi,\overline{\varphi}}\rangle-2\langle{\operatorname{grad}(\mu)\cdot V\cdot\varphi,\overline{\varphi}}\rangle,$
and for case 4, we get that
$\operatorname{Ric}(V,V)$ $=g(\operatorname{Ric}(V),V)=\langle{\varphi,\operatorname{Ric}(V)\cdot\overline{\varphi}}\rangle=-\langle{\operatorname{Ric}(V)\cdot\varphi,\overline{\varphi}}\rangle$
$=-\langle{2nV(\mu)\varphi+2\operatorname{grad}(\mu)\cdot V\cdot\varphi+4(n-1)\mu^{2}V\cdot\varphi+i(V\lrcorner\omega)\cdot\varphi,\overline{\varphi}}\rangle$
$=4(n-1)\mu^{2}g(V,V)-2\langle{nV(\mu)\varphi+\operatorname{grad}(\mu)\cdot V\cdot\varphi,\overline{\varphi}}\rangle$
$=4(n-1)\mu^{2}g(V,V)-2(nV(\mu)\langle{\varphi,\overline{\varphi}}\rangle-\langle{\operatorname{grad}(\mu)\cdot V\cdot\varphi,\overline{\varphi}}\rangle),$
which finishes the proof. ∎
Applying part 2 of Lemma 4, we can also compute the metric around any zero of $V$, when $\varphi$ is a special generalized Killing spinor.
Proposition 12.
Let $(M,g)$ be a time oriented semi-Riemannian $\text{spin}^{c}$ manifold of index $r$ and dimension $n$ carrying a special generalized Killing spinor $\varphi$ with Killing function $\mu$ such that any of the cases of Lemma 3 holds. Then the zeroes of $V$ are isolated. For each such zero $p\in M$, there exists a normal neighborhood $U\subset T_{p}M$, for which there are coordinates on $U\setminus C_{p}$, with $C_{p}$ denoting the light cone, where the metric takes the form $g_{(r,x)}=\epsilon dr^{2}+(\frac{\psi_{\epsilon}^{\prime}(r)}{\psi_{\epsilon}^{\prime\prime}(0)})^{2}\tilde{g}_{x}$, with $\epsilon=\text{sign}\hskip 2.84526ptg(V,V)$, and $\tilde{g}$ denotes the standard metric on $\{x\in\mathbb{R}^{n}\mid\langle{x,x}\rangle_{r,n-r}=\epsilon\}$. In these coordinates, we have that $V=\operatorname{grad}\psi$, where $\psi(r,x)=\psi_{\epsilon}(r)$. Furthermore, these coordinates extend to a conformally flat metric on all of $U$.
Proof.
This follows by Lemma 3 and Part 2 of Lemma 4. ∎
Remark 13.
In case the Killing function is a non-zero constant, i.e. the spinor is a non-parallel Killing spinor, then we have globally that $V$ is a gradient vector field $V=\operatorname{grad}\psi$. A straightforward case by case computation of $\operatorname{grad}\beta$ yields that the function $\Psi$ can, up to addition by a constant, be computed as
Case 1 Case 2 Case 3 Case 4
$\frac{i^{r+1}\beta}{2\mu}$ $\frac{i^{r}\beta}{2\mu}$ $\frac{\beta}{2\mu}$ $\frac{\beta}{2\mu}.$
In case $E$ or $\tilde{E}$ is complete, we get the following improvement of Proposition 11.
Proposition 14.
Let $(M,g)$ be a connected time oriented semi-Riemannian $\text{spin}^{c}$ manifold of index $r$ and dimension $n$ carrying a generalized Killing spinor $\varphi$, with $|V|$ nowhere vanishing, such that any of the cases of Lemma 3 holds. For any leaf $L_{p}$, the Weingarten map $W^{\tilde{E}}$ is as in (11). If $E$ is complete, then for any leaf $L_{p}$ we get that the map $\Phi^{L_{p}}:(\mathbb{R}\times L_{p},\frac{dt^{2}}{g(V,V)}+h_{t})\to(M,g)$, where $h_{t}=\Phi_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$, is a normal, semi-Riemannian covering. If instead $\tilde{E}$ is complete and $\varphi$ is special with $A=\mu\operatorname{Id}$, then for any leaf $L_{p}$, the map $\tilde{\Phi}^{L_{p}}:(\mathbb{R}\times L_{p},\epsilon\hskip 0.73975ptdt^{2}+f^{2}g|_{L_{p}})\to(M,g)$, where $f$ is as in (12), is a normal, semi-Riemannian covering. If in addition the function $\mu\beta$ is not constantly zero and either the function $\mu\beta$ or the function
  • $4(n-1)j^{2}\mu^{2}g(V,V)+2\overline{j}i^{r+1}(nV(\mu)\langle{\varphi,\varphi}\rangle+\langle{\varphi,\operatorname{grad}(\mu)\cdot V\cdot\varphi}\rangle)$ in Case 1 and Case 2
  • $4(n-1)j^{2}\mu^{2}g(V,V)-2nV(\mu)\langle{\varphi,\overline{\varphi}}\rangle-2\langle{\operatorname{grad}(\mu)\cdot V\cdot\varphi,\overline{\varphi}}\rangle$     in Case 3 and Case 4
is non-negative or non-positive, then the map $\tilde{\Phi}^{L_{p}}$ is an isometry for any leaf $L_{p}$.
Proof.
The computation of the Weingarten map is done in Proposition 11. The statements about the coverings follow from Lemma 3 and Proposition 9. The injectivity condition is a translation of the conditions of Corollary 10 and is in complete analogue to the proof of Proposition 11. ∎
The above results contain some geometric information of manifolds carrying generalized Killing spinors, but we are unable to draw conclusions about what type of spinor equation one can obtain on the fiber. However, when restricted to spin manifolds with Killing spinors, we have the following result due to Bohle [11, Thereom 6 &\& Theorem 7].
Theorem 15 (Bohle, 2003, [11]).
A time oriented semi-Riemannian warped product spin manifold of the form $(I\times F,\epsilon\hskip 0.73975ptdt^{2}+f^{2}g_{F})$ admits a Killing spinor with Killing number $\lambda\in\{\pm\frac{1}{2},0,\pm\frac{i}{2}\}$ if and only if the warping function satisfies $f^{\prime\prime}=-4\epsilon\lambda^{2}f$ and the fiber $F$ admits a Killing spinor to the Killing number $\pm\lambda_{F}$, where $\lambda_{F}^{2}=\lambda^{2}f^{2}+\epsilon\frac{1}{4}(f^{\prime})^{2}$.
Note that by rescaling the metric, one can always assume that Killing numbers is of the above modulus. The function computed in (12) satisfies the equation $f^{\prime\prime}=-4\epsilon(j\mu)^{2}f$ in case the special generalized Killing spinor of Proposition 11 is a Killing spinor with Killing number $\lambda\in\{\pm\frac{1}{2},0,\pm\frac{i}{2}\}$. Using Theorem 15, we improve Proposition 11 and Proposition 14 in the case of Killing spinors, by concluding that the fibers carries Killing spinors.
Theorem 16.
Let $(M,g)$ be a connected time oriented semi-Riemannian spin manifold carrying a Killing spinor $\varphi$ having Killing number $\lambda\in\{\pm\frac{1}{2},0,\pm\frac{i}{2}\}$, with $|V|$ nowhere vanishing such that any of the cases of Lemma 3 holds. Then for any uniform pair $(L_{p}^{\prime},I)$ for $\tilde{\Phi}$, we have that the map $\tilde{\Phi}^{L_{p}^{\prime}}:(I\times L_{p}^{\prime},\epsilon\hskip 0.73975ptdt^{2}+f^{2}g|_{L_{p}^{\prime}})\to(M,g)$, where $f$ is as in (12), is an isometry onto its image. The fiber $(L_{p}^{\prime},g|_{L_{p}^{\prime}})$ carries a Killing spinor with Killing number $\pm\lambda_{L_{p}^{\prime}}$, where $\lambda_{L_{p}^{\prime}}^{2}=\lambda^{2}f^{2}+\epsilon\frac{1}{4}(f^{\prime})^{2}$.
If $\tilde{E}$ is complete, then for any leaf $L_{p}$, the map $\tilde{\Phi}^{L_{p}}:(\mathbb{R}\times L_{p},\epsilon\hskip 0.73975ptdt^{2}+f^{2}g|_{L_{p}})\to(M,g)$, where $f$ is as in (12), is an isometry. The fiber $(L_{p},g|_{L_{p}})$ carries a Killing spinor with Killing number $\pm\lambda_{L_{p}}$ satisfying $\lambda_{L_{p}}^{2}=\lambda^{2}f^{2}+\epsilon\frac{1}{4}(f^{\prime})^{2}$.
Proof.
By proposition 11 we get that $\tilde{\Phi}^{L_{p}^{\prime}}$ is a local isometry. However, since the function $4(n-1)\lambda^{2}g(V,V)$, is either positive or negative, we have by the last condition of Proposition 11 that $\tilde{\Phi}^{L_{p}^{\prime}}$ is in fact an isometry. The statement regarding the Killing spinor on the fiber follows from Theorem 15. For the case when $\tilde{E}$ is complete, one argues identically, using Proposition 14. ∎
In the Lorentzian case, we have stronger results due to Bohle [11, Section 5] and Leitner [34, Theorem 5.3]. In the Riemannian case, the global statement is essentially the same as due to Baum in [7]. The author is not aware of similar local statements in the Riemannian case but is sure that they are known to experts. There are also stronger results for $r=2$, due to Shafiee and Bahrampour [38]. In case of parallel spinors we have stronger results in [16, 10, 39, 27].
5.2 The Riemannian classification
In this section, we specialize to the Riemannian setting, showing how one can generalize previous classifications made by Große and Nakad [23], and Leistner and Lischewski [33].
5.2.1 Type I imaginary generalized Killing spinors
We say that an imaginary generalized Killing spinor $\varphi$ on a Riemannian $\text{spin}^{c}$ manifold is of type I if the locally constant function $Q_{-}=\langle{\varphi,\varphi}\rangle^{2}-g(V_{\varphi},V_{\varphi})$ equals 0 everywhere. If $Q_{-}$ does not equal 0, then we say that $\varphi$ is of type II. By the Cauchy-Schwarz inequality, we always have that $Q_{-}$ is non-negative. It is a general fact that generalized Killing spinors have no zeroes. Indeed, a generalized Killing spinor is parallel with respect to the modified connection $\tilde{\nabla}^{S}_{X}\psi:=\nabla^{S}_{X}\psi-jA(X)\cdot\psi$ and thus if $\varphi$ had one zero, it would be zero throughout by $\tilde{\nabla}^{S}$-parallel transport. With this, we see in the Riemannian case that type I imaginary generalized Killing spinors are such that $|V_{\varphi}|\neq 0$. This is the first reason why we are able to classify type I imaginary generalized Killing spinors. As in Section 5.1, let $\Phi$ be the flow of $E:=\frac{V_{\varphi}}{|V_{\varphi}|^{2}}$, let $\tilde{\Phi}$ to be the flow of $\tilde{E}:=\frac{V_{\varphi}}{|V_{\varphi}|}$, and let $\mathcal{F}=\{L_{p}\}_{p\in M}$ be the foliation induced by $E^{\perp}$. Recall from Proposition 5 that $\Phi$ preserves the foliation $\mathcal{F}$ and if $\varphi$ is special, then $\tilde{\Phi}$ also preserves the foliation $\mathcal{F}$. The following lemma shows the second reason why we are able to classify type I imaginary generalized Killing spinors. The proof of the following lemma was pointed out to me by Jonathan Glöckle and is essentially the same computation as in section 1.3 of [22].
Lemma 17.
Given a spinor field $\varphi$ on a Riemannian $\text{spin}^{c}$ manifold such that $\langle{\varphi,\varphi}\rangle^{2}-g(V_{\varphi},V_{\varphi})=0$ everywhere, we also have that
$V_{\varphi}\cdot\varphi=i|V_{\varphi}|\varphi.$ (16)
Conversely, if $\varphi$ is an imaginary generalized Killing spinor which is not parallel on any component of $M$, and satisfies equation (16), then $\varphi$ is of type I.
Proof.
We compute that:
$\langle{V_{\varphi}\cdot\varphi-i|V_{\varphi}|\varphi,V_{\varphi}\cdot\varphi-i|V_{\varphi}|\varphi}\rangle$ $=\langle{V_{\varphi}\cdot\varphi,V_{\varphi}\cdot\varphi}\rangle-2\hskip 2.84526pt\text{Re}\langle{V_{\varphi}\cdot\varphi,i|V_{\varphi}|\varphi}\rangle+\langle{i|V_{\varphi}|\varphi,i|V_{\varphi}|\varphi}\rangle$
$=g(V_{\varphi},V_{\varphi})\langle{\varphi,\varphi}\rangle-2i|V_{\varphi}|\langle{\varphi,V_{\varphi}\cdot\varphi}\rangle+|V_{\varphi}|^{2}\langle{\varphi,\varphi}\rangle$
$=2g(V_{\varphi},V_{\varphi})\langle{\varphi,\varphi}\rangle-2|V_{\varphi}|g(V_{\varphi},V_{\varphi})$
$=2|V_{\varphi}|^{2}(\langle{\varphi,\varphi}\rangle-|V_{\varphi}|).$ (17)
First, assume that $\varphi$ is a type I imaginary generalized Killing spinor. Then (16) follows because both $g(\cdot,\cdot)$ and $\langle{\cdot,\cdot}\rangle$ are positive definite. Suppose now that $\varphi$ is an imaginary generalized Killing spinor which is not parallel on any component of $M$ such that $V_{\varphi}\cdot\varphi=i|V_{\varphi}|\varphi$. Then, for each point $x\in M$, we have by (5.2.1), that either $V_{\varphi}(x)=0$, or that $\langle{\varphi,\varphi}\rangle(x)-|V_{\varphi}|(x)=0$. Since the function $\langle{\varphi,\varphi}\rangle^{2}-g(V_{\varphi},V_{\varphi})$ is locally constant, it suffices to show that the set $N:=\{x\in M\mid V_{\varphi}(x)=0\}$ does not contain any component of $M$. Suppose for a contradiction that $N$ contains some component $M_{1}$ of $M$. By assumption, there is some point $x\in M_{1}$ and a vector $Y\in T_{x}M_{1}$ such that $A(Y)\neq 0$. Pick any local extension $\tilde{Y}$ of $Y$ to compute that
$0$ $=Y(g(V_{\varphi},A(\tilde{Y}))=Y(i\langle{\varphi,A(\tilde{Y})\cdot\varphi}\rangle)$
$=i(\langle{\nabla_{Y}\varphi,A(\tilde{Y})\cdot\varphi}\rangle+\langle{\varphi,\nabla_{Y}A(\tilde{Y})\cdot\varphi}\rangle+\langle{\varphi,A(\tilde{Y})\cdot\nabla_{Y}\varphi}\rangle)$
$=i(\langle{iA(Y)\cdot\varphi,A(Y)\cdot\varphi}\rangle+\langle{\varphi,A(Y)\cdot iA(Y)\cdot\varphi}\rangle)$
$=2\langle{\varphi,A(Y)\cdot A(Y)\cdot\varphi}\rangle=-2g(A(Y),A(Y))\langle{\varphi,\varphi}\rangle_{x},$
and since generalized Killing spinors have no zeroes, we conclude that $A(Y)=0$ which establishes the contradiction and completes the proof. ∎
Remark 18.
To see that the assumption that $\varphi$ is not allowed to be parallel on any component of $M$ in Lemma 17 is necessary, consider a complete and simply connected spin manifold $(M,g)$ that carries a parallel spinor and has irreducible holonomy. In this case, the Dirac current would have to vanish identically, because otherwise the manifold $(M,g)$ would carry a nowhere vanishing parallel vector field and the holonomy of $(M,g)$ would be reducible. Hence the Dirac current vanishes but $\langle{\varphi,\varphi}\rangle$ does not, and so $\varphi$ is of type II but satisfies equation (16) trivially.
Remark 19.
Baum introduced the terms type I and type II imaginary Killing spinors in [7]. It is shown in [7] and [8] that a type I imaginary Killing spinor on a Riemannian spin manifold satisfies (16). This proof can be adapted to imaginary generalized Killing spinors as well, but is very different from the above proof of Lemma 17. In [33], Leistner and Lischewski uses equation (16) directly.
In what follows, we will apply the $\text{spin}^{c}$ Gauß formula from [36, Section 3]. We follow the setup of that paper. Given a leaf $L_{p}\in\mathcal{F}$, we equip it with the induced $\text{spin}^{c}$ structure and we write $S_{L_{p}}$ to denote the induced spinor bundle on the hypersurface $L_{p}$. When the dimension $n$ of $M$ is odd, we have that $S_{L_{p}}\cong S|_{L_{p}}$ and when $n$ is even, we have that $S_{L_{p}}\cong S^{+}|_{L_{p}}$. Hence, when $n$ is odd, we can simply write $\psi|_{L_{p}}$ to denote the restriction of a section, but when $n$ is even we write $\psi|_{L_{p}}$, to mean $\psi^{+}|_{L_{p}}$. In this setup, the $S^{1}$-bundle associated with the $\text{spin}^{c}$ structure on $L_{p}$ is simply the restriction $P_{S^{1}}|_{L_{p}}$, where $P_{S^{1}}$ is the $S^{1}$-bundle associated to the $\text{spin}^{c}$ structure on $M$. Hence, given a connection 1-form $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$, we get induced connections $\nabla^{S}$ on $S$ and $\nabla^{S_{L_{p}}}$ on $S_{L_{p}}$. If we let $\bullet:T(L_{p})\otimes S_{L_{p}}\to S_{L_{p}}$ denote the Clifford multiplication on $S_{L_{p}}$, we have that $X\bullet\eta=\tilde{E}\cdot X\cdot\eta$, where on the right hand side, $\eta$ is either viewed as an element of $S|_{L_{p}}$ or $S^{+}|_{L_{p}}$, depending on if $n$ is odd or even.
Remark 20.
It is important to point out that the above identifications of spinor bundles and Clifford multiplication on hypersurfaces is not the only one. It boils down to the choice of inclusion $C\ell_{n-1}\to C\ell_{n}$ of Clifford algebras. The choice which is made in [36] is the inclusion $e_{j}\to e_{n}\cdot e_{j}$, where $\{e_{1},\ldots,e_{n}\}$ is the standard basis for $\mathbb{R}^{n}$ and thus also the set of standard generators for $C\ell_{n}$. Another choice that can be made is to choose the inclusion $e_{j}\to e_{j}$. The second choice would have worked equally well for the main results of this article. However, the below formula (18) would have had to been adjusted accordingly.
Translating Proposition 3.3 of [36], to our notation, we get for any spinor-field $\psi\in\Gamma(S)$ and any vector field $X\in\Gamma(T(L_{p}))$ that
$\nabla^{S_{L_{p}}}_{X}(\psi|_{S_{L_{p}}})=(\nabla^{S}_{X}\psi)|_{L_{p}}+\frac{1}{2}W^{\tilde{E}}(X)\bullet\psi|_{L_{p}}.$ (18)
With this, we use (11) to compute for any type I imaginary generalized Killing spinor $\varphi$ on a Riemannian $\text{spin}^{c}$ manifold, that the spinor $\varphi|_{L_{p}}$, satisfies
$\nabla^{S_{L_{p}}}_{X}\varphi|_{L_{p}}$ $=(\nabla^{S}_{X}\varphi)|_{L_{p}}+\frac{1}{2}W^{\tilde{E}}(X)\bullet\varphi|_{L_{p}}$
$=(iA(X)\cdot\varphi)|_{L_{p}}+\frac{1}{2}(\tilde{E}\cdot W^{\tilde{E}}(X)\cdot\varphi)|_{L_{p}}$
$=(iA(X)\cdot\varphi)|_{L_{p}}+(\tilde{E}\cdot(A(X)-g(A(X),\tilde{E})\tilde{E})\cdot\varphi)|_{L_{p}}$
$=(iA(X)\cdot\varphi)|_{L_{p}}-((A(X)-g(A(X),\tilde{E})\tilde{E})\cdot\tilde{E}\cdot\varphi)|_{L_{p}}$
$=(iA(X)\cdot\varphi)|_{L_{p}}-(A(X)\cdot\overset{=i\varphi}{\overbrace{\tilde{E}\cdot\varphi}}+g(A(X),\tilde{E})\varphi)|_{L_{p}}$
$=-g(A(X),\tilde{E})\varphi|_{L_{p}}.$ (19)
[…]
$\frac{X(\langle{\varphi,\varphi}\rangle)}{2\langle{\varphi,\varphi}\rangle}\varphi|_{L_{p}}=-g(A(X),\tilde{E})\varphi|_{L_{p}}=\nabla^{S_{L_{p}}}_{X}\varphi|_{L_{p}}.$ (20)
Lemma 21.
Let $\varphi$ be a type I imaginary generalized Killing spinor on a Riemannian $\text{spin}^{c}$ manifold. Then the spinor $\langle{\varphi,\varphi}\rangle^{-\frac{1}{2}}\varphi|_{L_{p}}$ is $\nabla^{S_{L_{p}}}$-parallel, for any leaf $L_{p}\in\mathcal{F}$.
Proof.
We compute that
$\nabla^{S_{L_{p}}}_{X}\langle{\varphi,\varphi}\rangle^{-\frac{1}{2}}\varphi|_{L_{p}}$ $=X(\langle{\varphi,\varphi}\rangle^{-\frac{1}{2}})\varphi|_{L_{p}}+\langle{\varphi,\varphi}\rangle^{-\frac{1}{2}}\nabla^{S_{L_{p}}}_{X}\varphi|_{L_{p}}$
[…]
and complete the proof. ∎
Remark 22.
By (5.2.1), we see that the condition for a type I imaginary generalized Killing spinor to restrict to a $\nabla^{S_{L_{p}}}$-parallel spinor on $L_{p}$ is that $g(X,A(\tilde{E}))=0$, for all $X\in T(L_{p})$. In case $\varphi$ in Lemma 21 is special, then this condition is satisfied and so $\varphi$ restricts to a $\nabla^{S_{L_{p}}}$-parallel spinor on $L_{p}$. This is the parallel spinor found in the proof of Theorem 4.1 of [23].
We can now state and prove our main theorems.
Theorem 23.
Let $(M,g)$ be a Riemannian $\text{spin}^{c}$ manifold carrying a type I imaginary generalized Killing spinor $\varphi$.
  1. a)
    For any uniform pair $(L_{p}^{\prime},I)$ for $\Phi$, the map $\Phi^{L_{p}^{\prime}}:(I\times L_{p}^{\prime},\frac{dt^{2}}{g(V_{\varphi},V_{\varphi})}+h_{t})\to(M,g)$ is a local isometry, where $h_{t}:=\Phi_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$ is a family of metrics on $L_{p}^{\prime}$, such that $(L_{p}^{\prime},h_{t})$ carries a parallel spinor for each $t$.
Suppose in addition that $\varphi$ is special with Killing function $i\mu$.
  1. b)
    For any uniform pair $(L_{p}^{\prime},I)$ for $\tilde{\Phi}$, the map $\tilde{\Phi}^{L_{p}^{\prime}}:(I\times L_{p}^{\prime},dt^{2}+e^{-4\int_{0}^{t}\mu ds}g|_{L_{p}})\to(M,g)$ is a local isometry and $\varphi$ restricts to a parallel spinor on $(L_{p}^{\prime},g|_{L_{p}^{\prime}})$. If $\mu$ is not identically zero on the image of $\tilde{\Phi}^{L_{p}^{\prime}}$ and is either non-negative or non-positive on the image of $\tilde{\Phi}^{L_{p}^{\prime}}$, then the local isometry $\tilde{\Phi}^{L_{p}^{\prime}}$ is an isometry onto its image.
Proof.
Everything except the last sentence follows directly from Proposition 11, Lemma 21 and Remark 22. To prove the last sentence, take a leaf $\Lambda\in\mathcal{F}$ that have non-trivial intersection with the image of $\tilde{\Phi}^{L_{p}^{\prime}}$ and such that $\mu|_{\Lambda}\neq 0$. Since the flow of $\tilde{\Phi}$ preserves the foliation, there exists a $t_{0}\in I$ such that $\tilde{\Phi}(t_{0},L_{p}^{\prime})\subset\Lambda$. The pair $(\tilde{\Phi}(t_{0},L_{p}^{\prime}),I-t_{0})$, where $(a,b)-t_{0}:=(a-t_{0},b-t_{0})$, is uniform for $\tilde{\Phi}$ and satisfies the first injectivity condition of Proposition 11. Therefore, the map $\Phi|_{(I-t_{0})\times\tilde{\Phi}(t_{0},L_{p}^{\prime})}$ is injective and since $\tilde{\Phi}$ preserves the foliation $\mathcal{F}$, the map $\tilde{\Phi}^{L_{p}^{\prime}}$ is also injective. Hence $\tilde{\Phi}^{L_{p}^{\prime}}$ is an isometry onto its image. ∎
Assuming completeness of $E$ or $\tilde{E}$, we get the following global analogue of the above theorem.
Theorem 24.
Let $(M,g)$ be a connected Riemannian $\text{spin}^{c}$ manifold carrying a type I imaginary generalized Killing spinor.
  1. a)
    If $E$ is complete, then for any leaf $L_{p}$, we have that $\Phi^{L_{p}}:(\mathbb{R}\times L_{p},\frac{dt^{2}}{g(V_{\varphi},V_{\varphi})}+h_{t})\to(M,g)$ is a normal, Riemannian covering, where $h_{t}=\Phi_{t}^{*}(g|_{L_{\Phi_{t}(p)}})$ is a family of metrics on $L_{p}$ such that for each $t$, the manifold $(L_{p},h_{t})$ carries a parallel spinor.
Suppose in addition that $\varphi$ is special with Killing function $i\mu$.
  1. b)
    If $\tilde{E}$ is complete, then for any leaf $L_{p}$, we have that $\tilde{\Phi}^{L_{p}}:(\mathbb{R}\times L_{p},dt^{2}+e^{-4\int_{0}^{t}\mu ds}g|_{L_{p}})\to(M,g)$ is a normal, Riemannian covering and $\varphi$ restricts to a parallel spinor on $(L_{p},g|_{L_{p}})$. If $\mu$ is not identically zero and is either non-negative or non-positive, then the covering $\tilde{\Phi}^{L_{p}}$ is an isometry.
Proof.
Everything except the last sentence follows from Proposition 14, Lemma 21, and Remark 22. To prove the last statement, one appeals to the injectivity condition of Proposition 14 and argues as in the proof of Theorem 23. ∎
Theorem 24 generalizes Theorem 4.1 and Corollary 4.2 of [23]. In particular, it extends Baum’s classification of type I imaginary Killing spinors on spin manifolds [7, Theorem 3] to the $\text{spin}^{c}$ setting. Theorem 24 a) also generalizes a part of Theorem 4 in [33].
5.2.2 Type II imaginary generalized Killing spinors
Recall that an imaginary generalized Killing spinor on a Riemannian $\text{spin}^{c}$ manifold is said to be of type II if the constant $Q_{-}=\langle{\varphi,\varphi}\rangle^{2}-g(V_{\varphi},V_{\varphi})$ is positive. In case the type II imaginary generalized Killing spinor is a Killing spinor, then Große and Nakad [23, Proposition 4.3] showed that the manifold is isometric to hyperbolic space, provided it is complete. It was shown by Rademacher [37] that on spin manifolds, all type II special imaginary generalized Killing spinors have constant Killing function, i.e. they are Killing spinors. Grosse and Nakad [23, Section 4.2.2] extends this to the $\text{spin}^{c}$ case for dimension at least $3$ and provide a counterexample in dimension $2$.
To clarify matters in the case when the generalized Killing spinor need not be special, we want to consider hypersurfaces of Lorentzian manifolds. The following identifications are described in [2] and [5] in the spin case and in [36] for the $\text{spin}^{c}$ setting. We give a quick summary similar to the above discussion on Riemannian hypersurfaces of Riemannian $\text{spin}^{c}$ manifolds. Let $(\overline{M},\overline{g})$ be a time oriented Lorentzian $\text{spin}^{c}$ manifold with spinor bundle $S$ and let $(M,g)$ be a spacelike hypersurface with unit normal $N$. We give $(M,g)$ the induced $\text{spin}^{c}$ structure and write $S_{M}$ for the corresponding spinor bundle. When $\dim{\overline{M}}$ is odd, we have that $S_{M}\cong S|_{M}$ and when $\dim{\overline{M}}$ is even, we have that $S_{M}\cong S^{+}|_{M}$. In this setup, the $S^{1}$-bundle associated with the $\text{spin}^{c}$ structure on $M$ is simply the restriction $P_{S^{1}}|_{M}$, where $P_{S^{1}}$ is the $S^{1}$-bundle associated to the $\text{spin}^{c}$ structure on $\overline{M}$. Hence, given a connection $A\in\Omega^{1}(P_{S^{1}},i\mathbb{R})$, we get induced connections $\nabla^{S}$ on $S$ and $\nabla^{S_{M}}$ on $S_{M}$. Clifford multiplication $\bullet$ on $S_{M}$ is related to Clifford multiplication $\cdot$ on $S$ via $X\bullet\psi=iN\cdot X\cdot\psi$. Correspondingly, the connection $\nabla^{S}$ relates to the connection $\nabla^{S_{M}}$, via
$(\nabla^{S}_{X}\psi)_{M}=\nabla^{S_{M}}_{X}(\psi|_{M})-\frac{i}{2}W^{N}(X)\bullet(\psi|_{M}).$ (21)
The positive definite hermitian bundle metric $(\cdot,\cdot)$ in $S_{M}$ from structures on the spinor bundle is given explicitly by $(\varphi,\psi):=\langle{N\cdot\varphi,\psi}\rangle$, where $\langle{\cdot,\cdot}\rangle$ is the scalar product in $S$. With this setup, we show that when generalized Killing spinors are defined in this generality, type II imaginary generalized Killing spinors exist in abundance.
Theorem 25.
Let $(\overline{M},\overline{g})$ be a time oriented Lorentzian $\text{spin}^{c}$ manifold carrying a parallel spinor $\varphi$. Then, for any spacelike hypersurface in $(\overline{M},\overline{g})$ with unit normal $N$, $\varphi$ restricts to an imaginary generalized Killing spinor with Killing endomorphism given by $\frac{1}{2}$ times the Weingarten map of that hypersurface. If the Dirac current $V_{\varphi}$, with respect to $\varphi$ on $(\overline{M},\overline{g})$ is timelike, then the restricted spinor is of type II, and if it is lightlike, the restricted spinor is of type I.
Proof.
By (21), we have that $\varphi$ restricted to any hypersurface is an imaginary generalized Killing spinor $\phi$ with Killing endomorphism $\frac{1}{2}W^{N}$. Let $U_{\phi}$ denote the Dirac current on the hypersurface with respect to this restricted spinor. For any vector field $X$ tangent to the hypersurface, we compute that
$\overline{g}(U_{\phi},X)$ $=i(\phi,X\bullet\phi)=i\langle{N\cdot\varphi,iN\cdot X\cdot\varphi}\rangle$
$=\langle{N\cdot N\cdot\varphi,X\cdot\varphi}\rangle=\langle{\varphi,X\cdot\varphi}\rangle$
$=-\overline{g}(V_{\varphi},X).$
Further, we see that
$\overline{g}(V_{\varphi},N)=-\langle{\varphi,N\cdot\varphi}\rangle=-\langle{N\cdot\varphi,\varphi}\rangle=-(\phi,\phi),$
and thus, along the hypersurface, we have that $V_{\varphi}=-U_{\phi}+(\phi,\phi)N$. If $V_{\varphi}$ is timelike, we get that
$0>\overline{g}(V_{\varphi},V_{\varphi})=\overline{g}(U_{\phi},U_{\phi})+(\phi,\phi)^{2}\overline{g}(N,N)=\overline{g}(U_{\phi},U_{\phi})-(\phi,\phi)^{2},$
which precisely means that $\varphi$ restricts to a type II imaginary generalized Killing spinor on the hypersurface. Similarly, if $V_{\varphi}$ is lightlike, the restricted spinor will be of type I. ∎
Lorentzian $\text{spin}^{c}$ manifolds carrying a parallel spinor such that the Dirac current is timelike, exists, and were studied in [28]. The second part of the above statement is essentially known and is a motivation for studying type I imaginary generalized Killing spinors; see for example, [33].